2 conducting thin concentric shells of radii r and 3r .outer shell carries charge q . inner shell is nuetral. Find charge that will flow from inner shell to earth after the switch is closed

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
At the point when S1 is shut 
Let charge q1 originates from the earth on external shell 
Vouter=0 
(Q2r+q12r)=0 [qi is charge on external circle after i times ] 
q1=−Q 
At the point when S2 is shut 
k(q|ir−Q2r)=0[qi| is charge on inward circle after i times] 
q|i=Q2 
Presently again when switch S1 is shut 
k(Q2×2r+q22r)=0 q2=Q 
Presently when S1 opened and S2 shut 
k(q|2r−Q22r)=0 
Continuing correspondingly we get after n times 
qn=−Q2n−1 and q|n=Q2n 
Potential contrast between shells after n times 
VAB=kq|n[1r−12r] 
VAB=kQ2n×2=kQ2n+1r.
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)